Instrument Log 

50 MHz WP C1 Lamont, OK Profiler Trailer

Date

Operator

Comments

1/9/07

PAD

Two brackets on SW speaker unit that is outside of the antennae field have worn through where tie down cables attach.At 2010 GMT, Mark Klassen and I replaced three plates on speaker frame where cables attach and adjusted all tie down cables.RTS at 2110 GMT.

5/5/06

PAD

Transmitter has not been operating for some time.Need to shut down instrument.At 1521 GMT, shut down computer, transmitter, amplifiers, and UPS for instrument.RTS at 1536 GMT.

5/4/06

Cam

Instrument mentor requested the instrument be shut down until the new tube is installed in the transmitter.Stopped program at 1850 GMT.

4/13/06

DKB

Powered off just the transmitter, pulled the VTX unit, and performed continuity checks for the heater.Found no problems.Still suspect the tube has failed.Notified the vendor and powered up the transmitter at 1618 GMT..

4/4/06

DKB

Installed the new computer motherboard and repaired VTX unit.Powered up the system and set the computer parameters to link with the SDS and set the date/ time.The RASS tested OK at 2100 GMT.The transmitter still has no power out and the wind data is bad.Left running at this time.

Fwd peak - 0, Fwd Ave - 0, Rev peak - 0, RF in - OK

LT volt - 48, STG1 current - 0.6A, STG2 current - 0.2A

HT volt - 8.2Kv, HT current - 0

No interlocks, fans/ fuses OK, HT/ Controller/ heater ON.Notified vendor and mentor.

1/3/06

PAD

Changed monthly operating parameters to January at 1628 GMT.

10/31/05

PAD

Last successful SDS collection was at 0922 GMT.At 1923 GMT, computer shows:Press spacebar NOW to invoke Hardware Profile/ Last Good Menu.Pressed spacebar and computer was not responsive.Power cycled computer.Program restarted correctly.RTS at 1938 GMT.

8/25/05

PAD

Last successful collection was on 8/24/05 at 2225 GMT.Computer display is black with blue box that shows symbols associated with monitor adjustments.Power light is flashing on computer and hard drive is inactive.Computer is not responsive.Power cycled computer at 1438 GMT.Program restarted correctly.After reboot, computer clock shows 0103.Instrument started RASS cycle.Reset clock at 1453 GMT. RTS at 1453 GMT.

8/9/05

PAD

SDS is not collecting data from instrument computer.Rebooted computer at 1412 GMT.RTS at 1427 GMT.††††

7/6/05

PAD

Last successful collection was at 1120 GMT.Rebooted computer at 1518 GMT.RTS at 1533 GMT.

6/7/05

MR

The new speakers for both SE speaker enclosures have arrived and need to be installed.Installed all 4 speakers in the enclosures.Tested the polarity of the wires and they tested OK.Powered up the speakers at 2055.The sounding at 2100 indicated that all speakers where working.Turned the power down to about 1/4 of the total power and will allow the new speakers to break in at this power until Friday.RTS at 2130 GMT.

5/13/05

MR/DN

The SW speakers where requested to be shut off.Shut off the speakers at 2115 GMT.

5/11/05

PAD

During PM rounds, observed transmitter control breaker in tripped position.Reset control breaker at 1454 GMT.RTS at 1509 GMT.

4/20/05

PAD

Speakers and box need to be removed from north RASS speaker to troubleshoot ongoing problem of speaker failure.At 1405 GMT, Mike Rainwater, Dan Nelson and I shut down 50 RWP and removed north RASS speakers and box.Restarted system at 1417 GMT.RTS at 1420 GMT.

4/8/05

MR

Green Country weed control is on site and needs to enter the antenna field.Shut the system down at 1305 GMT.Green country sprayed for weeds.They finished up at 1545 GMT.Powered up the system at 1550 GMT.RTS at 1600 GMT.

4/8/05

MR

The north RASS speakers have failed. Removed the speakers and will send them in for repair.Due to lack of spares, the north RASS amp will be left off.RTS at 1330 GMT.

4/6/05

PAD

At 1500 GMT, observed SW speaker was not operating.At 1820 GMT, Mike Rainwater and I removed and replaced SW RASS speakers.One speaker was frozen and the other tested 1 ohm.After installation, set amplifier output for speaker break - in period.RTS at 1920 GMT.

4/5/05

PAD

SW speaker not operating at 1600 GMT.Replaced fuse checked amplifier output power at 2030 GMT.RTS at 2045 GMT.

4/1/05

PAD

SW RASS speaker is not operating.Removed and replaced fuse for SW speaker.RTS at 1617 GMT.

3/29/05

MR

The instrument mentor wants the north speakers reinstalled and powered on.Installed and powered up the north speaker enclosure.RTS at 2130 GMT.

2/18/05

MR

The 50 RASS speakers and enclosures are ready to be reinstalled.Reinstalled the speaker enclosures and speakers.Tightened all the antenna wires in the field.Powered the system up at 2050 GMT.System came up and the speakers seemed to work.Powered the speakers up to the half way point and will allow the system to operate at this level until next week.RTS at 2215 GMT.

2/17/05

MR

The speaker boxes need to be rewired so that the speakers will run in series instead of parallel.Rewired the speaker enclosures.These enclosures will be reinstalled later this week.RTS at 1500 GMT.

2/15/05

PAD

Notified by facilities department 50 RWP can be started.At 1717 GMT, restarted transmitter, computer, receiver/ modulator and interface.Started LAP XM program.Event log shows DSP would not started.Acquire Data Failed.Found J3 cable from interface to computer was not seated properly.Reseated connector and program now running correctly.RTS at 1757 GMT.

2/15/05

PAD

50 RWP needs to be shut down so facilities department can perform electrical upgrades to profiler trailer. Stopped program at 1522 GMT.Shut down computer, transmitter, and UPS for receiver/ modulator and interface.Notified facilities.RTS at 1537 GMT.

1/19/05

PAD

At 1557 GMT, display shows dwell display and is updating. Computer clock is correct.Mouse is not responsive.Cannot bring up Lap XM.Control, Alt., Del. will bring up menu but cannot move cursor to select option.Called windows administrator.Tim was able to remotely access computer and reboot.Program restarted correctly and computer can be controlled locally.RTS at 1612 GMT.

1/14/05

MR

The south west inside and outside speakers need to be removed and sent back for repairs. Powered down the system at 1830 GMT.Removed the speakers.Powered up the RF part of the system at 2015 GMT.RTS at 2030 GMT.

12/28/04

PAD

Last successful collection was at 1022 GMT.At 1504 computer clock was current and time on lapxm display was updating.Clock on local dwell display shows 11:25:39 GMT.Stopped program and rebooted computer.Made sure program restarted correctly.RTS at 1519 GMT.

12/21/04

PAD

Observed Control switch on Atrad transmitter was off.Shut down and restarted transmitter as per written instruction.RTS at 1600 GMT.

12/16/04

DN

Since relocation of the SE RASS on 12/3/2004, two of the three RASS speakers have failed (SW1 & N). Review of the Crown audio amplifier manual seemed to indicate that we should be operating the amplifiers in Bridge-Mono, rather than Parallel-Mono mode, based on the measured load impedance of 8.5 ohms for the SW2 RASS speaker.The SW2 (relocated SE RASS) audio amplifier was configured for Bridge-Mono and set to minimum gain. During the 2300 RASS cycle, the amplifier gain was slowly increased in stepwise fashion, while monitoring the RASS output. During confirmation testing, the remaining RASS speaker failed at 50% of maximum gain (~700W). The gain in Parallel-Mono mode is normally set at 80-85% of maximum and the speaker is rated at a minimum of 800W continuous power. The amplifier has since been configured back to Parallel-Mono mode, and a recommendation of installing protective inline speaker fuses made, as strongly suggested in the Crown manual.

12/13/04

PAD

SW RASS speakers located in antenna field are not working.At 2111 GMT, Mike Rainwater and I checked SW RASS speakers.Speakers tested open or burnt out.Shut down amplifier.Waiting on replacement speakers.RTS at 1610 GMT.

12/6/04

MR

The instrument was started at 1450 GMT but it was observed that not all of the new speakers where operating.T/S found that the north speakers where not working.Further T/S found that the new speakers that where installed had blown.Removed the speakers but due to lack of spares, this channel of the system will remain off.Checked the remaining two amps for clipping of the the signal and adjusted properly so that there was no clipping.Allowed the system to run on its own at 1800 and did not observe any problems.RTS at 1815 GMT.

11/15/04

PAD

Computer department notified SDS system has not collected data.Computer has a C:\ defrag box open.Stopped program and restarted computer at 1836 GMT.RTS at 1850 GMT.

11/9/04

PAD

50 RWP last successful collection was at 1122 GMT.Collection log showsftpcopy:got unwanted answer to `RETR // SpecMom/H04314a.SPC':550/ SpecMom/H04314a.SPC:The process cannot access the file because it is being used by another process.Restarted computer at 1549 GMT.Made sure program restarted correctly.RTS at 1603 GMT.

8/5/04

PAD

Last successful collection was at 1120 GMT.Rebooted computer to reestablish communications at 1514 GMT.RTS at 1529 GMT.

7/28/04

MR

The site power will be going down and this instrument needs to be shut down.Shut the instrument down between and 1500 GMT.RTS at 1530 GMT.

7/27/04

MR

The North rass speakers seemed to not be working. T/S found the speaker wire had vibrated loose on one of the speakers and caused the other speaker to be overloaded and burned out.Shut the program and RF amp down at 1948 GMT.Removed both speakers and will ship off for repair.Could not replace the speakers at this time due to lack of spares.Restarted the RF amp and program at 2025 GMT.RTS at 2035 GMT.

7/15/04

PAD

During troubleshooting amplifier problems on 7/14/04, amp for SW RASS speakers was left turned off.Powered up amplifier at 1457 GMT.SW speaker was operating during 1500 GMT RASS cycle.RTS at 1512 GMT.

7/14/04

CAM

The north RASS amplifierís ODEP circuitry was activating about midway through the RASS cycle and shutting off the output for that speaker.T/S during the 1800 GMT RASS period found that the amplifierís internal fan had failed and was causing the amplifier to overheat.Due to failed speakers in the southeast RASS, the southeast amplifier is not being used at this time.Moved the north speaker to the southeast amplifier and confirmed good operation during the 1900 GMT RASS period.RTS at 1910 GMT with the north and southwest RASS units operational.The southeast RASS remains down pending in-house replacement of the amplifier fan and vendor repair of the speakers.

7/13/04

CAM/PAD

The north and southeast RASS speakers were not working.Shut the system down at 1828 GMT to allow entrance to antenna field.Checked speaker wires for shorts and continuity.No problems found.Measured the north speakerís impedance at the external terminal and found that it was open.Removed the two speakers from the north enclosure and found that the push button terminal was broken on one of the speakerís positive terminal, which allowed the wire to vibrate out and lose connection, but otherwise the speaker measured ok.The other speakerís coil measured open and had failed.R/R both speakers in the enclosure and reinstalled the north RASS.††† Measured the southeast speakerís impedance at the external terminal and found that it too was open.Removed the speakers from the southeast enclosure and measured both speakers individually.Both coils were open.There are no more spares available to replace at this time.†† Will return the 4 bad speakers to the vendor for repair.The system will be run with only the north and southwest RASS speakers operational until spares become available.The southeast amplifier will be left powered off.Restarted the system at 2106 GMT.The system came on line during the last few seconds of the RASS period, at which time the north and southwest units were confirmed operational.RTS at 2110 GMT.

6/30/04

MR

The power to the site was lost at 1330 GMT.While power was out, the system was physically turned off so a controlled power up could be performed.Power was restored to the site at 1455.Powered up the system at 1607 GMT.System came up and rebooted ok.RTS at 1622 GMT.

6/9/04

PAD

50RWP computer displayed error:Internet Explorer cannot open the internet site http://sgp_50/scripts/Lapxm_Isapi_Home.dll?Mfcl SAPI Command=Start and Config=June 5 dfcf.cfg&x=28&y=4.Operation timed out.LapXM display showed graph but dwell box was empty and control panel was not displayed.Rebooted computer at 1541 GMT.Made sure program restarted correctly.RTS at 1551 GMT.

4/29/04

KT

The parameter and configuration monthly header was changed from April to May.

4/29/04

KT

One of the speaker cables that comes from underground had been chewed by ground animals.At 1505 GMT the exposed wiring appeared in good condition so I took electrical tape and fixed area where the wire was exposed.RTS at 1520 GMT.

4/29/04

KT

Green Country weed control was on site to spray the 50 RWP antenna field.At 1456 GMT stopped the program and powered down the 50 RWP.At 1657 GMT Green Country had finished spraying the antenna field.At 1700 GMT powered up the 50 RWP.RTS at 1705 GMT.

4/23/04

PAD

Computer clock was 1 min. 13 sec. fast.Adjusted computer clock at 1523 GMT.RTS at 1538 GMT.

4/20/04

MR

The SDS has lost communications with the instrument.Rebooted the 50 RWP computer.Communications was restored.RTS at 1525 GMT.

3/29/04

PAD

50RWP program had locked up on 3/28/04 at 1911 GMT.Stopped program at 1535 GMT.Restarted computer at 1539 GMT.Inadvertently unloaded configuration file.Error displayed:Unable to load configuration file.Reloaded Mar Configuration file.Restarted lapxm program at 1555 GMT.RTS at 1557 GMT.

3/11/04

PAD

During PM rounds, computer clock was 1 min. 10 sec. fast.Reset computer clock at 1549 GMT.RTS at 1604 GMT.

3/5/04

PAD

50 RWP display did not show Lapxm control panel.Error message displayed:Microsoft internet explorer cannot open the Internet site http://sgp_50Scripts/ Lapxm_Isapi_Home.dll?mfclSAPICommand=Start&Config=Mar5dfcf.cfg&x=23&y=4.Stopped and restarted program at 1539 GMT.Made sure program restarted correctly.RTS at 1554 GMT.

3/1/04

JEM

The 50RWP config file needs to be changed to the current month.Stopped data acquisition @1628 GMT.Updated to the March config file and edit the file header the reflect the month of March.Started Data acquisition @1630 GMT.Data appears normal.RTS @1643 GMT.

2/26/04

PAD

Corner of tarp covering north speaker box is not secured.††† Stopped 50 RWP program at 2028 GMT.†† Secured tarp to frame using tie-straps.Check other tarps and replaced tie-straps as needed.Started program at 2038.Made sure program restarted correctly.RTS at 2043 GMT.

1/15/04

PAD

Instrument had not collected since 0320 GMT.Rebooted computer at 1446 GMT, and made sure program started normally.Instrument collected data.RTS at 1501 GMT.

12/3/03

CAM

The 50 RWP computer clock was 2 minutes fast.Set 50 RWP computer clock to SDS time at 1952 GMT.

11/6/03

MR

The 50 RWP speakers have returned and need to be placed back into the system.Replaced the speakers.RTS at 2145.

10/24/03

PAD

50RWP has not collected data since 0920 GMT.Computer program was running and had reasonable data.Rebooted computer at 1612 GMT.Made sure computer program was running.Computer collected data at 1620 GMT.RTS at 1627 GMT.

10/16/03

PAD

Last successful collection was at 1020 GMT.Computer was running and time was current.Graph was displayed but did not update as normal.Rebooted computer at 1356 GMT.Made sure program started and was normal.RTS at 1406 GMT.

9/23/03

KC

Shut down Transmitter and computer at 2020 GMT. Pulled the speakers from the North and Southeast Drum. Power cycled Transmitter and computer at 2055 GMT.RTS at 2300 GMT.

9/23/03

KC

Speaker Drum tarps were loose. Shut down Transmitter and computer at 1508 GMT. Tightened down tarps, power cycled Transmitter and computer at 1522 GMT. RTS at 1523 GMT.

9/22/03

PAD

UPS for 50 RWP was not putting out any voltage.Removed UPS model BC 900 Lan and replaced with (2) *UPSís.Power cycled computer and checked that program was running RTS at 1951 GMT.*New - LI 660VB 54071*New - LI 660VB 29659

9/22/03

KC

50 RWP computerwas not running. Checked reset of power strip, found UPS battery backup to be the problem. The UPS was not putting out any power, and the 50 receiver, monitor , and computer power strip were plug into the UPS.Plugged the Amplifier into the power strip, along with monitor. Plugged power strip directly into wall socket at 1505 GMT. RTS at 1515 GMT.

9/12/03

CAM

SDS instrument status page indicated that the 50 RWP collections had been failing since 1220 GMT.†† Daily rounds check earlier today found the computer was at the login prompt, but once logged in the program was observed running normally.Rebooted the 50 RWP computer at 1801 GMT.RTS status O.K. at 1805 GMT.

8/21/03

CAM

SDS instrument status page indicated that no files had been received from the 50RWP since 1020 GMT today.Check of computer found the program was running normally.Rebooted the computer at 1811 GMT.System RTS at 1814 GMT.The following collection was successful at 1820 GMT and all missing data was recovered.

8/8/03

JEM

LapXM is not running and there are approximately 150 illegal operation errors on the screen.Closed all the error boxes and restarted the 50RWP computer @1338 GMT.At 1344 GMT, the 50RWP computer was back online.Started LapXM and loaded the August configuration file @1345 GMT.At 1347 GMT, the 50RWP was operating normally.RTS @1350 GMT.

8/7/03

KT

The 50 RWP has not been heard rassing.At 1610 GMT stopped the program to troubleshoot.At 1649 GMT inspections of the cables and connections were complete.No problems were found.At 1657 GMT started the program.The program started to run but kept on stopping and starting by itself.At 1708 GMT called Rich Coulter and notified mentor of situation.Mentor asked to have the input voltage checked for the RASS board in the computer.Voltage was 0.Instrument mentor believes possible RASS board bad after receiving the computer back from repairs.Toubleshooting still in progress.Waiting to hear from mentor for further instructions.RTS at 1740 GMT.

8/6/03

MR

The instrument mentor wants the computer card checked and reseated and the connections to the amps checked due to the instrument not outputting any audio signal.Checked the connections to the amp and they checked good.Checked and reseated the computer card and it checked good.Notified the mentor.

7/29/03

TG

Received and installed the 50RWP computer today at 1930 GMT. Powered up computer and amps at 1945 GMT. System returned operational status O.K.

7/17/03

TG

Received and installed the 50 RWP computer yesterday.When the computer booted up it halted at the BIOS screen, with the A Drive light staying on constant. The computer would not allow use of the keyboard.Contacted Gerold at Vaisala today to see if he might have used a different keyboard during setup or if there might have been some changes made to the configuration of the computer. He said that the computer worked O.K. there with the same keyboard/mouse configuration that we have. After explaining how the computer is acting, Gerold said that the problem might be with the CPU / Motherboard. Gerold requested opening the case to inspect the interior.Opened the case and observed that the CPU heat sink / Fan had come off and was lying against the CPU / motherboard and the CD Rom Drive mounting chassis.Reinstalled the CPU heat sink / Fan and then booted up the computer. The computer still would not get past the initial Boot screen. Gerold said that the CPU Heat sink / fan could have become dislodged during shipment.Pulled the computer to ship back to Vaisala for repairs since it is still covered under warranty.

7/7/03

TG

After working with Gerold Lipford from Vaisala regarding the new computer boot failures, he suggested that the computer be returned for warranty repair.Removed the 50 RWP Computer and packed it for shipment to Vaisala.

6/26/03

JEM

The 50RWP computer was found with the following error.Windows NT could not start because the following file is missing or corrupt:\WINNT\SYSTEM32\CONFIG\SYSTEM Hardware Profile/Last Known Good Menu.Tried restarting the computer several times, but always received the same error.Notified Tim Grove and Rich Coulter about the problem.No further action taken at this time.Did not RTS at this time.

4/17/03

JEM

When upgrading the 50RWP the mentor observed that the southwest speakers were not functioning.Powered the system down @1825 GMT.Disconnected the 50RWP speakers @1835 GMT.Removed both speakers from the speaker box and replaced from 1840 GMT to 1940 GMT.Connected new speakers and replaced the tarp @1945 GMT.Powered the system back on @1955 GMT.At 2000 GMT the 50RWP began to RASS as expected and all speakers were working properly.RTS @2005 GMT.

3/5/03

JEM

The cards from the 50RWP computer need to be shipped to Vaisala so a new computer can be built for the 50 RWP.Stopped the program @1545 GMT.Powered off all amplifiers for the 50RWP.Disconnected all cables from the computer and delivered it to the lab for the cards to be removed and prepared to be shipped.RTS @1600 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.Powerwas restored to the Central Facility at approximately 1945 GMT.RTS @2000 GMT.

1/8/03

TG

Technician stopped the pop4 program at 1655 GMT and ran defragr.exe on C Drive at 1657 GMT. Technician also deleted D00, D01, and D02 log files from C:\radar\data directory. Computer now has 161 Meg free hard drive space. System was rebooted at 1717 GMT and RTS at 1719 GMT. Status O.K.

1/8/03

JEM

The 50RWP has not sent any files since 1/7/03 @2310 GMT.The 50RWP computer was found to be locked up during "data transfer after consensus".Powered the computer off @1505 GMT.Powered the computer back on @1506 GMT.At 1512 GMT the 50RWP appeared to be functioning normally and was displaying 10dB.RTS @1515 GMT.

1/3/03

JEM

The 50RWP locked up during its data transfer after consensus.Powered the 50RWP computer off @1516 GMT.Powered the computer on again @1517 GMT.At 1525 GMT, the 50RWP appeared to be operating normally displaying 11dB.RTS @1530 GMT.

12/16/02

JEM

The 50RWP has lost its network connection, and has become stuck trying to establish a connection.Powered the 50RWP computer off @1517 GMT.Powered the computer back on @1518 GMT.The 50 RWP computer was back online and operating normally @1525 GMT.RTS @1530 GMT.

11/12/02

MR

The instrument is asking for a new password.The old password for the system has been changed.Called the computer department and Trent changed the BIOS so the new password would be accepted.RTS at 1550.

10/24/02

KT

SDS indicated that the 50 RWP was not sending any files.At the time of inspection the computer for the 50 RWP was found to be locked up.†† At 1452 GMT rebooted the computer.RTS at 1506 GMT.

9/13/02

JEM

The 50 RWP has locked up while waiting for the 1300 RASS.Shut down the computer @1405 GMT.Powered the computer on @1407 GMT.Status returned to normal and the data appeared reasonable @1420 GMT.RTS @1420 GMT.

8/14/02

CAM

SDS comlog reports indicated that no files had been received from the 50RWP since 1410 GMT and the computer can not be pinged. T/S found the program was not running and the computer displayed an error connecting to network drive f:.Power cycled the computer at 1746 GMT.The program restarted and system RTS status O.K. at 1749 GMT.The following collection was successful at 1810 GMT.

8/14/02

MR

The SDS has lost communications with the instrument.Rebooted the computer at 1430 GMT.RTS at 1445 GMT.

8/1/02

JEM

There is no AC power to the instrument.A breaker for the outlets had tripped.Powered off all amplifiers and the computer and reset the breaker.Powered on all amplifiers and computer @1400 GMT.System was operational and data appeared normal @1410 GMT.RTS @1410 GMT.

7/29/02

MR

While starting the dehumidifier the breaker for the computer was tripped.Reset the breaker and rebooted the computer.RTS at 1455 GMT.

7/18/02

CT

According to the SGP instrument status page the instrument has not collected files since 7/17/02 at 2210 GMT. The instrument was rebooted at 1345 GMT.Status OK. RTS at 1400 GMT.

7/12/02

KT

The North Rass speaker cover needed to be secured back to the frame.At 1520 GMT stopped the program and reattached the cover to the frame.At 1529 GMT started the program.RTS at 1535 GMT.

6/17/02

CT

The computer needed to be rebooted. Rebooted computer at 1349 GMT. RTS at 1404 GMT.

6/14/02

MR

Power restored.System operational.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.The instrument is shut down at this time.

5/10/02

KT

The SDS reported that the instrument was not collecting data. The computer was power cycled at 1415 GMT. RTS at 1430 GMT.

5/9/02

KT

Due to severe storms the 50 RWP lost power.At 1350 GMT power was restored but still no power was going to the 50 RWP.At 1427 GMT observed that the #2 breaker was tripped.At 1430 GMT reset the breaker and still no response.At 1433 GMT powered off all computers and amplifiers and reset the breaker.After the breaker was reset all computers and amplifiers were powered on one at a time.RTS at 1437 GMT.

5/7/02

CT

The noise level could not be determined and the radar monitor was not displaying the cloud cover. The computer was power cycled at 1401 GMT. RTS at 1416 GMT.

3/11/02

MR

The power to the site was lost over the weekend.The instrument was found to be off and the circuit breaker was found to be tripped.Reset the circuit breaker and powered up the amps.Powered up the computer.The system came up and started taking data at 1533 GMT.RTS at 1546 GMT.

3/4/02

CT

Instrument was down due to the breaker being tripped. Instrument was powered on at 1444 GMT. RTS at 1459 GMT.

2/19/02

CT

The computer was displaying the password prompt. The password was entered at 1523 GMT. RTS at 1538 GMT.

2/14/2002

TMG

Computer will not load TCPIP on bootup and fails to connect to network. Technician R/R 16 bit network card and loaded new card drivers. Computer would not work with AUI port. Technician configured network card to use Twisted Pair and TCPIP would load. Computer can only be rebooted by power OFF/ON and not Ctrl-Alt-Del or network card will not work properly.Rebooted computer and RTS 2/14/2002 at 1700 GMT.

2/11/02

MR

The AC power to the site was lost on 1/30/02 due to local ice storm.Powered up the instrument at 2045 GMT.The computer restarted but the DOS could not find the network.Rebooted the system again with the same results.Shut the system down and Reseated the network card.Restarted the system again with the same results.Allowed the system to continue to boot without finding the network and they system was able to come up and start taking data.Notified the computer department.The network card is suspected of going bad.Will replace as soon as one is located or a suitable replacement is found.Computer department recommended shutting the system down until then.Shut the system down at 2130.

12/21/2001

CJ

The amplifier filter needed to be cleaned. Cleaned the filter at 1446 GMT. Reinstalled at 1501 GMT. RTS at 1501 GMT.

12/18/2001

TMG

Technician stopped data acquisition program at 1935 GMT to configure system for NFS change over. System RTS at 2129 GMT.

11/23/01

CAM

SDS comlog reports indicated that 50 RWP data had not been ingested since 2122 GMT on 11/21.Rebooted the 50 RWP computer at 1456 GMT to reestablish the network connection.RTS at 1458 GMT.

11/20/01

MR

The power to the site will be shut off for about 1 hour due to a site water leak repair.Instrument was shut down from 1500 to 1600 GMT.RTS at 1615 GMT.

11/08/2001

CAM

Technician edited autoexec.bat to set computer clock with new time server address. RTS at 1625 GMT.

11/05/01

CJ

Due to severe weather the program had terminated. Rebooted computer at 1435 GMT. RTS at 1450 GMT.

11/01/2001

CJ

The tarps that are on the 50 RWP speaker frames needed to be tightened due to high winds. Secured the tarps at 1600 GMT. RTS at 1630 GMT.

10/29/01

CJ

Time was found to be incorrect on the screen at 1542 GMT. Reset to correct time at 1542 GMT. RTS at 1557 GMT.

8/20/01

MR

The power was lost to the site on 8/15/01 from 0030 GMT to 0400 GMT.Due to the instruments being down the link to the SDS was lost and needs to be reset. Rebooted the instrument at 1510 GMT to reset the link.RTS at 1525 GMT.

7/2/01

JEM

Computer displays for both 50 and 915 RWP instruments were totally blank at 1335 GMT. Performed soft reboot (cont/alt/del) at 1340 GMT. Both displays were functional at 1350 GMT.

5/30/01

MR

Instruments and computers were down. Made sure all instruments were powered down. Reset circuit breaker #2 at 1436 GMT. Also power cycled UPS. Instruments came back up at 1437, and all are operational.

4/25/01

JEM

The 50 RWP needs to be shut down to allow the sprayers access to the field so that can spray to kill the weeds and grass that is growing there. Stopped the program on the computer @1531 GMT. Shut down the computer @1533 GMT. Shut down all amplifiers and turned off all the breakers that go to the 50RWP @1538 GMT. The Sprayers entered the area @1545 GMT to spray. Turned on the breakers, amplifiers, and the computer @1855 GMT. The 50 RWP was observed functioning properly @1900 GMT.

4/10/01

TG/CM

The mount directory needed to be changed for the new collection system. Shut the system down at 2046 GMT and changed the mount directory to /files1/5.7/data/collection/sgp/sgp50rwpC1.00. Rebooted the computer and RTS at 2102 GMT.

3/13/01

MR

The system is continually RASSing and will not stop. Rebooted the system at 1630 GMT. The system came back up and upon start up of the RASS program the system started RASSing again and would not stop. Called Rich Coulter and consulted with him. He thinks that maybe the parameter file was loaded in on 3/11/01 and has not been changed back. Rich loaded the correct parameter file into the system and the system started working. RTS at 1730 GMT.

3/6/01

MR

The RF transmitter is not working. Shut the transmitter down at 1518 GMT. Removed the cover and checked for loose connections. Did not observe anything obvious. Replaced the cover and powered the transmitter back on. The RF meter was showing good output and the data on the display was good. RTS at 1548 GMT.

2/24/01

TG

Tim Grove had to come in on 2/24/01 to reboot the computer systems after the site power was out from 1125 to 1253 GMT. The main circuit breaker for the computers and amps had tripped when site power was restored. Upon resetting the circuit breaker for the 915 computer, the breaker tripped immediately. Unplugged the 915 amp. This allowed the computer and the 50 amps to be powered up at 1945 GMT. T/S will continue for the 915 amp on 2/26/01.

1/26/01

MR

The computer had the error "General failure Reading drive F: " Rebooted the computer at 1508 GMT. System came back up and started running status OK at 1509 GMT. RTS at 1523 GMT.

11/2/2000

MR

Some of the tie wraps for the RASS speaker covers have come loose. Replaced the tie wraps and checked for loose ones.

1/26/01

MR

The computer had the error "General failure Reading drive F: ".Rebooted the computer at 1508 GMT. System came back up and started running status OK at 1509 GMT. RTS at 1523 GMT.

9/6/00

MR

Stopped the program and rebooted the computer at 1600 GMT due to it not sending files. System came up and started collecting at 1602 GMT. RTS at 1615 GMT.

6/30/2000

CM/GS

Instrument mentor requested that the Peavey audio amplifiers be replaced with new Crown Com-Tech 1610 model amplifiers. Powered the 50 MHz Profiler down at 1455 GMT. R/R the *SW, **N, and ***SE audio amplifiers. Initially the amplifiers were configured for bridged mono mode (4/8 ohm) as requested by mentor. System power was restored at 1720 GMT. Checked on RASS status at 1800 GMT and found that the amplifiers IOC (Input/Output Comparator) status light lit after about 6 minutes of operation. This indicated that the output was being distorted. The ODEP (Output Device Emulation Protection) light then went out shortly after which indicated that the output impedance was too low. The speaker impedance is roughly 4 ohms and bridged mono mode requires a minimum 8 ohm load. Powered the amplifiers off at 1810 GMT and reconfigured all three for parallel mono mode. This is the recommended configuration for loads of 4 ohms or less. Restored amplifiers power at 1840 GMT. Verified RASS operation at 1900 GMT and set the input attenuation level to 1.0 dB for maximum sound level without distortion. The amplifiers ran the full RASS period with no distortion. RTS status O.K. at 1910 GMT. *Old - 6542586 *New - 128902 **Old - 8812035 **New - 128913 ***Old - 8700798 ***New - 128914

6/30/2000

CM/GS

Instrument mentor requested that a new upgraded DSP RASS board be installed in the 50 MHz profiler computer. Powered the 50 MHz computer off at 1455 GMT. Removed the old *DSP RASS board and installed the upgraded board. Restored power after additionally installing new amplifiers and RTS status O.K. at 1840 GMT. Verified RASS operation at 1900 GMT. RTS status O.K. *Old - 05104 *New - 001327

6/12/00

KT

The computer was locked up and needed to be rebooted. At 1325 GMT rebooted the computer. At 1328 GMT computer came on line. RTS at 1340 GMT.

4/24/00

KT

The NORTH speaker cover needed to be reconnected to the speaker frame. At 1810 GMT performed repairs. RTS at 1902 GMT.

4/3/00

KT

The computer needed to be rebooted due to network problems. At 1818 GMT rebooted the computer. RTS at 1822 GMT.

3/30/00

KT

At 1824 GMT the 50 RWP computer was locked up. At 1827 GMT rebooted the instrument computer. At 1931 GMT computer came on line. RTS at 1935 GMT.

1/26/00

MR

Mentor wants all the data files from 1/1/00-1/7/00 downloaded into the F: drive. Downloaded all the *.CNS and the *.SPC files to the F: drive.

1/25/00

MR

Brian Ermold requested that we check and see if there is data on the system from Jan 1-7. Stopped the program at 1523 GMT. Checked the DATA directory and found the files that where requested. Called Rich Coulter and notified him. Also emailed Brian Ermold and notified him. Rebooted the system at 1525. System started to run at 1528 status O.K.

01/07/00

TMG

Stopped Radar program at 2101 GMT and copied a new transfer.exe file to C:\Radar. Rebooted computer at 2135 GMT and RTS.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will need to be shut down. Shutdown 50 RWP at 1445 GMT, restarted 50 RWP at 1600 GMT. RTS at 1600 GMT.

12/14/99

TMG

Stopped Radar Program at 2015 GMT to perform Y2K software upgrade as per Tim Martin. Performed Y2K upgrade and rebooted computer at 2032 GMT. Computer displayed software version is POP V4.36.01. System RTS at 2035 GMT. Status O.K.

8/9/99

MR

Shut instrument down due to site construction at 1320 GMT.

7/13/99

MR

Started the reinstallation of the final RF according to the diagram in the installation manual on 7/9/99. After installation the Fault Temp light would immediately come on when the program started to run. Called the Kalumus company (manufacturers of the amp) and consulted with one of their service techs. He suggested that the 50 MHz computer was not working properly or that maybe something had vibrated apart during shipping of the amp and that I should remove the cover and inspect the amp. On 7/12/99 I removed the cover on the amp and examined but found nothing unusual. Reinstalled the cover and the amp. Called Rich Coulter and described the problem to him. He suggested calling Mr. Newshafer at Radian. Called Mr. Newshafer and he recommended checking the duty cycle of the gate signal coming from the computer. Checked the duty cycle and found it was at approximately 15% which is below the recommended 20% for the amp to function properly. On 7/13/99 I checked the input and output of the T/R switch and noticed the output of the switch was a negative pulsed signal. This was suspected of causing the final RF amp to go into the fault/temp mode. Rechecked the wiring of the system and found that the wiring diagram in the 50 RWP installation manual was not correct when compared to the diagram in the RF final amp installation book. Changed the TTL GATE signal cable from the positive pulse input to the negative pulse input on the back of the RF amp. This corrected the problem. RTS status O.K. at 1930 GMT.

6/25/99

SS

Removed the final amplifier at 1745 GMT, to be shipped for repair. RTS 1815 GMT.

6/23/99

CAM

The RF amplifier power meter showed 0 Watts for both forward and reflected power and the data did not appear reasonable. Stopped the 50 RWP program and rebooted the computer at 1755 GMT. The program restarted at 1758 GMT and the power meter then showed power but only about half as much as normal. The data then appeared reasonable. The power began slowly dropping and returned to 0 by 1803 GMT. Rebooted the computer again at 1817 GMT. RTS at 1819 GMT and once again power was indicated but only about 1/4 of the scale and was varying. The power dropped to 0 again at 1825 GMT. At 1827 GMT the power meter again showed about 1/2 the normal power but returned to 0 at 1829 GMT and remained there. The data however still appears reasonable. RTS as is at 1840 GMT. The mentor has been notified of the situation.

6/23/99

CAM

Daily status check found that the RF amplifier power meter indicates 0 Watts for both forward and reflected power. The data also does not appear reasonable. All normal status lights are lit on the amplifier. Checked RF amplifier fuse to be OK. Power cycled the amplifier with no change in status. Left voice mail for mentor detailing the problem.

6/16/99

TMG

Stopped program at 1836 GMT and edited autoexec.bat for the computer to write to cfhost instead of nfhost. Rebooted system at RTS at 2007 GMT.

6/16/99

TMG

Stopped program at 1537 GMT and edited autoexec.bat to keep the computer from writing to nfshost during upgrade to cfhost. Rebooted system at 1535 GMT and RTS.

5/26/99

MR

PM. Morning rounds showed the 50 was rassing but was paused. Reloaded the parameter file "STD_P~00" and rebooted. Checked on the 50 at 2:00 and it seemed to be running O.K.

5/25/99

CAM

The 50 RWP was hung up in RASS only mode. Instrument mentor had attempted to load a new parameter file but the T1 line went down and he was no longer able to gain access to the system. The RASS was sounding continuously but the program was on pause. Copied G:\radar.nus to C:\radar\radar.par and rebooted computer at 2020 GMT. This parameter file was RASS only as well. Stopped the program and copied the May5 parameter file to c:\radar\radar.par. Rebooted computer at 2030 GMT. Program RTS in normal mode at 2032 GMT. Mentor will check out the problem with the other parameter files when the T1 line is restored.

5/25/99

MR

PM. SDS lost comms with the instrument. Rebooted and let Mr. Coulter reload the parameter file. RTS at 1615. Status O.K.

5/13/99

MW

Further troubleshooting revealed that the new speaker that was installed on 5/12/99 was wired in backwards. The speaker leads were reversed and the RASS sounded normally. The 50 RWP was out of service from 1630-1821 GMT while troubleshooting.

5/12/99

MW

The 50 RWP has been operating with the SW RASS only since two of the RASS amplifiers burned up on 4/28/99 during continuous operation testing. The SW RASS was sounding at below normal volume. Powered the 50 RWP down at 1713 GMT. Found one of the speakers in the SW RASS enclosure was open circuited. R/R the blown speaker. Checked all RASS speakers with an ohmmeter. The North speakers ohmed out at 6 ohms each with a total line impedance of 4.5 ohms (speakers plus wire). The Southeast speakers ohmed out at 6 ohms each with a total line impedance of 4.5 ohms. The Southwest speakers ohmed out at 6-7 ohms with a total line impedance of 4.7-5.0 ohms. Installed two new *amplifiers. Both the Southeast and North RASS units sounded normally but the Southwest RASS was sounding at a very low volume. RTS at 2100 GMT with only the North and Southeast RASS operational. Will continue T/S tomorrow. *New - CS-800X SN 08812035 *New - CS-800 SN 00-05551727

5/10/99

JEM

The SDS was unable to communicate with the 915 RWP. Rebooted the 50 RWP. RTS @1455 GMT.††††

4/29/99

MR

Rebooted the computer due to the NFS host being rebooted at 1513. RTS at 1520.

4/28/99

MR

Copied the Config.sys and the autoexec.bat files to the G: drive for Jess Smith. Removed the *IP 8.5C amp and the **CS800X amp due to failures during mentors RASS only test. *Old - 8872418 **Old - 006542586

4/26/99 -4/27/99

Mentor

Instrument mentor on site for maintenance and RASS evaluation. Two Peavey amplifiers failed during the maintenance.

4/23/99

MR

Power cycled the computer at 1340 due to no comms with the SDS.

4/20/99

CAM

Work order #3422 calls for operating the 50RWP in continuous RASS mode using only one amplifier and speaker for one hour on 4/20/99. This is to test the hardware durability for an upcoming IOP campaign. Rich Coulter has arranged to have the continuous RASS parameter file load automatically at 1600 GMT and then return to normal operations at the 1700 GMT RASS period. Contacted Rich at 1555 GMT to verify the new parameter file was set to go as planned. Powered off the north and southwest amplifiers. Reduced the gain of the remaining southeast amplifier by 3dB. Monitored the system to be sure that the RASS remained on continuously from 1600 to 1700 GMT. No problems were encountered during the test period. Powered the north and southwest amplifiers back on at 1700 GMT and increased the gain of the southeast amplifier back to its original setting. Verified that the original parameter file was reloaded and the system returned to RASS + winds mode at 1710 GMT. RTS status O.K.

4/12/99

MW

The SW RASS amplifier MAINS circuit was tripping whenever the RASS would sound. Powered the system down at 1822 GMT. One of the SW RASS speakers was shorted. Replaced the shorted speaker and powered back up. RTS at 1912 GMT.

4/9/99

CAM

Several of the tie wraps used to secure the tarps to the speaker enclosures were broken. Powered the 50 RWP down at 1952 GMT. Entered the antenna field and secured the tarps with tie wraps. Restored power and RTS at 2012 GMT.

1/22/99

CAM

Shutdown profile program and modified autoexec.bat to correct POP4 program at 1837 GMT. The following command was changed: idnet login oper password to idnet login new user new password. Edited c:\pctcp\pctcp.ini and changed: user = oper to user = new user. Saved configurations and RTS at 1846 GMT.

12/28/98

MW

The 50RWP system and two RASS speaker amplifiers had no power. Found circuit breakers number 2 and number 9 tripped. Reset both circuit breakers at 2130 GMT. RTS at 2133 GMT.

12/23/98

CAM

Network Administrator requested that the 50 RWP computer be rebooted following network (nfshost) maintenance. Rebooted 50 RWP computer at 2120 GMT. RTS at 2122 GMT.

12/02/98

TMG

Technician stopped radar program at 1531 GMT to install PCTCP for DOS software. Installed FTP software and edited the autoexec.bat file with the following additions:

Idrive

Idnet login oper (password)

Idmnt f: (NFS Host in directory path)

Idmnt g: (NFS Host out directory path)

Setclock ntphost address

Technician rebooted computer and copied all radar data files to NFS Host. After all files were transferred, technician deleted all data files from c:\radar\data directory. 50rwp-cf rebooted and RTS at 1558 GMT.

11/24/98

TMG

SDS NFS will not accept data files from 50rwp-cf computer, therefore the 50rwp-cf computer hard drive has 23.5 Meg of free space. Technician shutdown 50rwp RASS program at 2012 GMT and ftp'd all data files to NFS Host. After files were transferred, Technician deleted all files from c:\radar\data directory. 176.8 Meg free space is available on the hard drive. System rebooted and RTS at 2045 GMT.

11/17/98

TMG

SDS NFS will not accept data files from 50rwp-cf computer, therefore the 50rwp-cf computer hard drive has 49 Meg of free space. Technician shutdown 50rwp RASS program at 1522 GMT and ftp'd all data files to hawk ftp server for storage until SDS is ready to accept the files. After files were transferred, Technician deleted all files from c:\radar\data directory. 176.8 Meg free space is available on the hard drive. System rebooted and RTS at 1600 GMT.

11/10/98

TMG

SDS nfshost will not accept data files from 50rwp-cf computer. 50rwp-cf computer has 27 Meg of free hard drive space for file storage. Technician exited RASS program at 2015 GMT. Technician ftp'd all c:\radar\data files (151 Meg) into a data storage drive on ops2 g:\50rwp_data. Technician deleted all files from c:\radar\data allowing 176 Meg of free space on the hard drive. Rebooted computer and RTS at 2100 GMT.

10/29/98

TMG

Computer host settings needs to be changed to nfshost instead of cf10. Technician stopped program at 1710 GMT and edited c:\nfs\host file. Technician changed the host table from 198.124.99.10 cf10 to 198.124.99.12 nfshost.sgp.arm.gov. Technician then edited c:\nfs\drives.bat to the following:

net use f: cf10:/files5/data/raw/50rwpC1/in

changed to

net use f: nfshost:/files5/data/raw/50rwpC1/in

net use g: cf10:/files5/data/raw/50rwpC1/out

changed to

net use g: nfshost:/files5/data/raw/50rwpC1/out

Rebooted Computer at 1752 GMT and RTS. Status O.K.

10/13/98

MR

Rebooted the computer due to no files being sent to the SDS.

10/5/98

MW

The SE RASS amplifier mains had been tripped sometime over the weekend. Turned the amplifier power switch off. Pressed the mains switch on the back of the amplifier and the main circuit breaker tripped. R/R the *Peavey CS-800X amp at 1710 GMT. RTS at 1745 GMT. *Old - 08812035 *New - 00-06542586

9/24/98

TG

The mentor has requested that the system computer be checked for Y2K compatibility and defragment the Hard Drive. Technician stopped program at 1614 GMT. Technician created c:\defrag directory and copied a:\defragr.exe, a:\defragr.doc and a:\2000.exe to c:\defrag directory. Rebooted system with bootup disk to c: >. Ran chkdsk /f to fix the FAT (26 bad clusters). Ran 2000.exe to check for Y2K compatibility. Software indicated "This system correctly supports the Year 2000". Ran defragr.exe to defragment the Hard Drive. Rebooted computer and RTS at 1630 GMT. Status O.K.

8/26/98

MW

Found the N RASS IP 8.5 amplifier circuit protection breaker tripped at 1545 GMT. Backed the input sensitivity down to 9dB and reset the amplifier circuit protection breaker at 1545 GMT. The IP 8.5 amplifier cannot drive the RASS units with the same power as the CS800X amplifiers without tripping its circuit protection breaker. As a result, the N RASS is now operational but at a significantly reduced signal amplitude compared to the SE and SW RASS units.

8/25/98

MW

The SE RASS has been off-line since 7/29/98 due to a partially shorted speaker. Exited the "POP4.BAT" program at 1417 GMT. Shut down the 50RWP at 1420 GMT. R/R the partially shorted speaker in the SE RASS. Powered the 50RWP back on and rebooted the computer at 1545 GMT. Set the IP8.5 amplifier input sensitivity setting to 3dB (same as the other two amplifiers.) During the first RASS at 1700 GMT, the IP8.5 amplifier SPS LED started to flicker on approximately 6 minutes into the RASS. Eventually, the SPS LED stayed lit until the end of the RASS. Swapped the North and Southeast RASS lines so the IP 8.5 amp would be powering the N RASS during the next RASS period. The amp exhibited the same characteristics with the N RASS as it did with the SE RASS. Let the system run overnight to see if the amplifier circuit protection breaker would eventually trip.

7/29/98

MW

On 7/29/98, noticed the SE IP8.5 amplifiers SPS light coming on intermittently. Powered the amp off at 1409 GMT. Shut the entire system down at 1614 GMT and measured each speaker with an LCR meter at 120 Hz. Measured one speaker at 10 ohms at 120 Hz (which is normal) and the other at 6 ohms at 120 Hz. The speaker which measured 6 ohms at 120 Hz did not have near the output of the other speaker. Suspect that one of the speakers in the SE RASS is damaged/partially shorted and this is causing the amps protection circuitry to kick on. Left the SE RASS off line and powered the system back on at 1650 GMT. RTS with the North and Southwest RASS operational at 1653 GMT.

7/28/98

MW

On 7/28/98, measured the SW RASS speaker line at 3.7 ohms with an ohmmeter and 2.65 ohms at 120Hz with an LCR meter. Shut the 50RWP down at 1805 GMT and replaced a shorted speaker in the SW RASS. Powered the 50RWP back on at 2050 GMT.

7/27/98

MW

The 50RWP has been experiencing problems with the amplifiers clipping and sometimes burning up, especially on the SW and SE RASS amps. The amplifiers usually fail once one of the speakers fails (shorts). The RASS speakers on the 50RWP take a lot of abuse when the signal is applied because the signal is applied instantly at full power, instead of gradually increasing the power like the 915RWP RASS. On 7/27/98, a new *CS-800X amp was installed on the SW RASS. When the 1900 RASS began, the SW amp DDT LED's immediately came on. Powered the amp off to prevent damage. *New - 08812035

7/23/98

MW

The SW RASS Peavy CS 800 amplifier blew a fuse this morning during the 1500 GMT RASS. Ohmed out the SW speaker line at 4 ohms. Disconnected everything from the amp and replaced the fuse and turned the amp on. The fuse instantly blew again. Removed the *amplifier from service at 1300 GMT. No spare amplifier to replace. *Old - 00-05551727

7/20/98

MK

During daily PM checks it was discovered that the SW RASS speaker was barely audible. Disconnected speaker wire from amplifier and ohmed out the speakers, speakers measured 9.4 ohms. Powered the RF amplifier off at 1506 GMT. Entered the antenna field and removed the RASS unit for repair. Ohmed out the speakers, one of the speakers was open, the other tested ok. Replaced speaker with rebuilt spare. Installed the repaired RASS unit and powered RF amplifier back on at 1812 GMT.

7/13/98

MR

Powered up the instrument at 1622. System O.K.

7/10/98

CAM

The Profiler trailer air conditioner has failed and must be powered off until repairs can be made. This will result in extremely high temperatures inside the trailer. The 50 RWP was powered down at 1843 GMT to prevent damage to computer and electronics from high temperature.

6/30/98

MK

The southeast amplifier overload protector has been tripped. Reset the protector at 1629 GMT. Status OK.

6/29/98

MK

The 50 RWP data is not actively collected by the SDS since 6/26/98. Rebooted the 50 RWP computer at 1815 GMT. RTS at 1830 GMT.

6/18/98

SS

Daily check revealed that the south east RASS amplifier power indicator was not lit. Turned power off to amplifier and reset unit circuit breaker. Turned power on and power indicator remained lit. RTS at 1645 GMT.

6/16/98

MW

SDS operations reported no data files were being collected from the 50 RWP. Rebooted the system at 1937 GMT. RTS at 1950 GMT. SDS operations reports data collection has resumed after the reboot.

6/15/98

CM/SS

The SE RASS speaker was barely audible. Powered off the RF amplifier at 1935 GMT to enter the antenna field. Ohmed out the speakers and they checked out OK. RF amplifier power restored at 1958 GMT. Checked operation of speaker during 2000 GMT RASS. Increasing the gain of the amp had no effect. Connected the SE speakers to a spare amplifier but they still could not be heard. By this time the RASS interval had expired. Powered the SE amplifier off and will continue T/S tomorrow.

6/15/98

SS

SDS reports bad ingest. Assume a loss of network connection. Shut down and rebooted computer. RTS at 1750 GMT.

6/2/98

CAM/MK

Received two repaired IP8.5 amplifiers. The 50 RWP was running without a SE amplifier. Installed repaired *SE amplifier at 1850 GMT. Adjusted the gain to +7.5 dBV which was maximum gain without SPS circuitry being active. All amps operational as of the 1900 GMT RASS. *New - 08771910

6/2/98

MK

Due to a power failure on 6/1/98, the 50 RWP has lost its connection with the network. At 1713 GMT, rebooted the 50 RWP computer. Status OK. RTS at 1729 GMT.

5/26/98

MK

Due to a power failure on 5/25/98 at 0752 GMT, the 50RWP computer has lost its network connection. Shutdown and restarted the 50 RWP computer at 1713 GMT, status OK. RTS at 1730 GMT.

5/19/1998

SS

During daily checks found the north RASS amp power not lit. Turned off power switch and reset unit main breaker and turned power on. main breaker immediately tripped. Turned power off and reset breaker and disconnected audio input and turned power on. Main breaker immediately tripped. Disconnected input and output from amp, turned off power and reset breaker. Turned power switch on and breaker immediately tripped. Pulled *amp to be repaired. *Old - 06542586

5/18/98

CAM

The power meter on the RF Amplifier showed zero forward power being transmitted. The winds data spectrum looked jumbled to around 5 km and was constant at zero beyond that. Power cycled the RF Amplifier at 1850 GMT with no change in status. Pressed the ALC button several times and the forward power finally returned to normal at 1855 GMT. The data now looks good throughout the spectrum. RTS status O.K. at 1900 GMT.

5/15/98

SS

Power outage. Rebooted computer to put it back on network.

5/7/98

SS

System rebooted due to power failure. Rebooted system at 1518 GMT.

4/22/98

MW

The Northern RASS has not been operational since 1/2/98 when its amplifier failed. The amplifier was returned to the SGP CF yesterday after repairs were finished. Installed the Northern RASS *amplifier and verified operation 1610-1710 GMT. Amplifier model CS-800X. RTS at 1710 GMT. *New - 00-06542586

4/14/98

KT

Re-booted the 50 RWP computer due to NT Server not operating properly. RTS at 2014 GMT.

4/14/98

KT

Re-booted the 50 RWP computer due to NT Server not operating properly. RTS at 1520 GMT.

4/6/98

MW

The amplifier for the Northern RASS was found with no power. Checked circuit breaker and power at outlet. Power OK. Power cycled the amp's power switch with no response. Suspect the amp's power supply burned up. Removed the *amplifier from service at 1810 GMT. *Old - 08700798 model CS 800X.

1/2/98

GRS

Morning rounds found that the N and SE RASS amps internal breakers were tripped. They tried to reset the main reset for the SE amp and it reset until the 1900 RASS and then it tripped the main breaker shortly after it started RASSing. When the N amp was reset it sent a spark out of the back of the amp, tripped its breaker, and a small amount of smoke was seen. They left both the N and SE amps powered off for further T/S when time allowed. T/S found that the N RASS amp was dead. Tried to run the SE RASS amp with the N amp off during the 1900 RASS. SE amp still tripped after only running a couple of minutes. R/R the N amp. It ran O.K. during the 2000 GMT RASS. The SE amp still tripped after only running a few seconds. Reduced the db gain back to zero and it would still trip. Moved the SE speaker wires to the new N amp and the SE speakers worked fine. Removed the SE amp. RTS the 50RWP at 2011 GMT with out a SE amp. * SE amp removed - MDL# IP-8.5C SN# 08872418. * N amp removed - MDL# CS-800X SN# 0006542586. * New N amp - MDL# CS-800X SN# 08700798.

12/31/97

GRS

The morning rounds found that the SE and N RASS units were not as loud as they should be. Suspect bad speakers. Shut the system down to troubleshoot the RASS speakers. SE RASS speakers and speaker wires OHM out O.K. The North speakers OHM'd out one bad speaker. Swapped the SE and SW speaker wires. Powered up the system at 1721 GMT. The SE RASS speakers worked on the next RASS at 1800 but the SW now does not. Suspect a bad SE RASS amp. Shut the system down again at 1813 to R/R the SE RASS amp and to remove the North RASS speaker box. R/R the SE RASS amp. Swapped the SE and SW speaker wires back to where they belong. Powered up the system at 1828 with no N RASS speakers. Moved the speaker box to the spares building to R/R the bad speaker. While T/S the bad speaker it was found that a speaker wire had broken where it connects to the speaker. Stripped the wire back and placed the wire back in the spring loaded receptacle on the speaker. Performed a check of the speakers and box before reassembling the speaker box. Tests and looks O.K. Shut the system down at 1930 to place the North speaker back were it belongs. Placed speaker on its rack and tie rapped the North and SE canvas speaker box covers back in place. Powered the system up at 1945 GMT. The North speaker now works properly, but the SE speaker does not. The channel A decibel gain on the suspect amp was set to 9db. Adjusted the replacement amp to 9db when swapped out. The amp would not power the speaker at 9db. Adjusted the db gain slowly up to infinity with no change. Adjusted the db gain slowly down to zero. The RASS speaker worked only between the 3 to 6 db. The RASS speaker worked best at 4.5 db. Left the SE amp adjusted to 4.5 db. Will check out the suspect amp. It possibly works and just needed the db gain adjusted to 4.5 range. RTS the 50 RWP at 2110 GMT with all RASS components working. Amps Model # IP8.5C * Old - 08771910 * New - 08872418

12/8/97

CAM

After further investigation Rich Coulter found that the parameter files minimum range was 4C so on cold days all of the data points were out of range and therefore interpreted as invalid. Mentor remotely loaded the correct parameter file at 2015 GMT.

12/8/97

CAM

SDS TR indicates that the temperature profiles are still reporting 9999's. Spoke with Rich Coulter and after explaining the situation to him and how the data appeared on the monitor, Rich suspects the parameter file may be corrupted. During the previous two reboots the data on the monitor has been a perfectly straight vertical line at 4C. Rich will check into it and get back with us.

12/5/97

CAM

Temperature profiles are still reporting 9999ís even after the soft boot on 12/4. Performed a complete shut down of system at 1613 GMT. RTS at 1617 GMT. Will continue to monitor status.

12/4/97

CAM

SDS TR indicated that the temperature profiles are reporting all 9999ís. Rebooted computer using ctrl-alt-del at 2050 GMT. RTS at 2055 GMT. Will monitor status via SDS to see if problem is corrected.

11/26/97

CAM

SDS TR indicates that the temperature profiles are reporting mostly -999's. However, the data on the PC monitor looked good. The SE RASS unit did not appear to be as loud as the other two units. System was powered down at 1619 GMT to enter field and check the speakers of the SE RASS, as well as to perform a general reboot in an attempt to clear the -999 problem. Both speakers in the SE unit checked OK at 6 ohms. System RTS at 1638 GMT. SDS operator then indicated that the temperature profiles returned to normal following the reboot.

9/25/97

CM/MW

During daily PM checks it was discovered that the north RASS unit was not operational. The exact time of outage is not immediately known. Powered the RF amplifier off at 1715 GMT. Entered the antenna field and removed the RASS unit for repair. Powered RF amplifier back on at 1722 GMT. T/S found that both speakers in the unit were blown (measured infinite resistance). Replaced both speakers with rebuilt spares. Powered RF back off at 1847 GMT and installed the repaired RASS unit. Restored RF power at 1854 GMT and all RASS units operational as of the 1900 GMT sounding.

9/17/97

MW

The Northern most RASS does not sound as loud as it should. It definitely is not as loud as the other two RASS units. No corrective maintenance performed at this time.

9/3/97

CAM

Due to a Kay Electric transformer going down in Billings on 9/2, the site lost power at around 2008 GMT. Site power was restored at 0142 GMT on 9/3 but because the SDS had also shut down, the 50 RWP did not establish a network connection properly upon power up. As a result, no files had transferred to SDS since 1910 GMT on 9/2. Rebooted 50 RWP computer at 1936 GMT. The network connection still did not appear to be established properly. Rebooted computer again at 1945 GMT. At 2000 GMT 23 files were successfully transferred to SDS. System RTS status O.K. at 2006 GMT.

7/18/97

CAM

The input sensitivity of the SE amplifier had been turned to +12dBV to prevent clipping of the signal on 10/24/96. It was later determined that one of the speakers was going bad (measured 3.6 ohms) and had to be replaced. The input sensitivity was never adjusted back after the speaker was replaced. Both the N and SW amplifiers sensitivity are set to +4.5dBV. Adjusted the SE amplifier input sensitivity to +4.5dBV but again the SPS circuitry was active during RASS mode indicating that the signal was clipping. The sensitivity had to be backed off to a maximum of +7.5dBV before the clipping stopped. RTS at 1910 GMT.

7/16/97

GRS

The 50 was getting weeds growing back in the antenna field. Performed a graceful shutdown of the instrument at 1452 GMT to allow a company to spray weed killer on the vegetation. After spraying was completed, turned the breaker on the RF amplifier and power cycled the 50ís UPS to reboot the system at 2112 GMT. Started taking data at 2113. RTS at 2114 GMT.

7/14/97

CM/ZW

The north RASS speaker was not operational. T/S found that both speakers in the unit were bad (infinite resistance). Powered the system off at 1412 GMT and removed the north RASS unit to the storage building to R/R speakers. Powered the system back on at 1420 GMT without the north RASS. Removed the faulty speakers and replaced with spares. 50 RWP powered back down at 1840 GMT. Installed the repaired north RASS unit and system RTS at 1856 GMT. All RASS units operational as of the 1900 GMT sounding.

6/13/97

MK/ZW

Vegetation is growing inside the antenna array. RTS at 1414 GMT.

6/2/97

MR

Rebooted system at 1620 GMT due to SDS losing the communications address.

5/15/97

ZW

The speaker housing replacement tarps came in. Powered off the main RF amplifier at 1812 GMT. Removed the old tarps, and replaced them with new modified tarps. Powered on main RF amplifier at 1935 GMT. Status ok. RTS 1940 GMT.

5/12/97

ZW

The North speaker housing has a plastic tarp in place of the missing tarp. Due to strong wind the plastic tarp became loose. Powered off the main amplifier to the 50 WP at 1415 GMT. Secured the loose tarp and powered on the main amplifier at 1420 GMT. Status ok. RTS 1430 GMT.

4/17/97

ZW

The tarps on the back of the speaker housings are showing signs of wear and tear. The tarps are in need of replacement. At 1310 GMT the RF Amplifier was shut off so measurements could be taken of the tarps so replacement tarp could be made. The RF Amplifier was powered on at 1318 GMT. Status ok.

3/20/97

ZW/CM

From 2211 3/19/97 SDS reported all 9999ís. Observed that the RF Amplifier was without power and found the RF Amplifier breaker was tripped. Reset RF Amplifier breaker at 2015 GMT. Status ok. RTS 2030 GMT.

2/27/97

TG/CM

Bill Horstemke requests that site ops place a new transfer.exe file on the 50 MHz wp/rass from g: drive to correct problems with file naming. Shutdown program at 1822 GMT and copied c:\radar\transfer.exe to c:\radar\transfer.sav. Deleted the old c:\radar\transfer.exe file and copied the new transfer.exe file from g: drive to c:\radar\transfer.exe. Rebooted system and RTS at 1826 GMT.

2/25/97

TG/CM

Bill Horstemke requests that the new transfer.exe file that site ops placed on the system be removed and the saved transfer.sav file be restored to its original file due to unforeseen data naming problems. Shutdown program at 2040 GMT and copied c:\radar\transfer.exe to c:\radar\transfer.tmg. Copied c:\radar\transfer.sav to c:\radar\transfer.exe. Rebooted system and RTS at 2047 GMT.

2/21/97

TG

Shutdown program at 2013 GMT to install a new transfer.exe file. Renamed old c:\radar\transfer.exe file as c:\radar\transfer.sav and installed new transfer.exe to c:\radar. Rebooted system at 2016 GMT and RTS. Note: Will check transfer status 02/24/97 before installing at IFís.

2/12/97

CAM

No files being received by SDS since 0111 GMT on 2/11/97. SDS System Administrator advised to reboot at 1625 GMT. As of 1811 GMT still no files being received. SDS T/S to continue.

12/17/96

TMG

Exited program at 1601 GMT to modify software for transferring files. RTS at 1608 GMT.

12/05/96

CAM

Reinstalled north RASS unit. RF powered off from 1836 to 1850 GMT. All RASS units operational as of 1900 GMT sounding.

11/20/96

CAM

North speaker inop at 1508 GMT. Amp is on but no output. SE amp (IP8.5C) SPS light is flashing. Will T/S.

11/20/96

CAM

T/S indicated both speakers in the north RASS unit were blown. System powered off at 1832 GMT. Removed north RASS unit to storage facility and removed speakers to be sent in for repair. System RTS at 2012 GMT with only 2 RASS units operational.

11/8/96

CAM

Found North Rass amplifier circuit breaker tripped. Reset and tested OK at 1606 GMT. DDT light is active intermittently.

10/28/96

CAM

Daily check found the 220 VAC breaker tripped over the weekend due to T-storms. Reset and tested OK at 1446 GMT.

10/25/96

CAM

Powered off system and R/R SE speaker. RTS at 1946 GMT.

10/24/96

CAM

Powered off system to T/S SE speakers. One speaker measures 3.6 ohms (low). Will replace tomorrow. RTS 1954 GMT.

10/21/96

CAM

Found SE amplifier circuit breaker tripped. T/S indicated that there may be a problem with the speakers. Will T/S when weather permits.

10/18/96

CAM

Found SE and North Rass amplifier inop. T/S found the trailer AC outlets tripped. Reset breaker and tested OK at 1600 GMT.

10/16/96

CAM

T/S SE Rass amplifier shutting off after 5 minutes. * R/R and tested OK at 2000 GMT. * old - 06315588 * New - 06542586

10/15/96

CAM

* Received and installed repaired amplifier. * New - 06315588

10/2/96

GRS

Found trailer very hot. AC control unit had lost itís program. Reset OK. The 220 V AC breaker had tripped. Reset and tested OK at 1448 GMT.

9/30/96

CAM

* Removed faulty amplifier. * Old - 06542586

9/25/96

CAM

Received repaired Rass amplifier, * installed and tested OK at 1935 GMT. RTS all 3 Rass at 2000 GMT. * new - 06542586

9/16/96

GRS

Daily check found 220 V AC breaker tripped. Reset and tested OK at 1440 GMT.

9/13/96

CAM

Received new Amplifier (8.5), * installed on the North speakers, and tested OK at 1500 GMT. * new - 08771910

9/12/96

CAM

T/S found that the amplifier would only run for a few minutes, shut off for 1 minute, and then continue for the rest of the Rass period (3 min.) Contacted vendor who informed us to send the amplifier back with the operating parameters that we are currently using. *Removed amplifier. Still only operating with 1 Rass amplifier. * old - 06315588

9/11/96

CAM

Installed repaired amplifier and tested. It would only run for 5 min. before shutting itself off. Powered off and will continue to T/S tomorrow. * new - 06315588

9/4/96

CAM

North Rass amplifier circuit breaker tripped. During T/S smoke stared coming from amp. * Removed amplifier and sent out for repair. This left only one Rass amplifier on line. * old - 06542586

8/27/96

CAM

Powered up system and tested OK.

8/22/96

CAM

Powered off system due to trailer AC unit failed. System off line until repairs can be made.

8/19/96

CAM

East amplifier circuit breaker had tripped. Reset the breaker at 1502 GMT. Amplifier sparked and smoked but breaker did not trip. Powered amplifier off and * removed for repair at 1530 GMT. * old - 06315588

8/13/96

CAM/GRS

The SW RASS assembly sounded as if one of the two speakers was not working. T/S found that one of the speakers was blown. Powered the system off at 1740 GMT. Removed the blown speaker and replaced it with a new spare. System RTS at 2014 GMT.

8/08/96

CAM

1400 GMT. VSWR will not reset.

8/07/96

GRS

1400 GMT. VSWR light flashing. Tried to reset but would not reset.

8/05/96

CAM

1400 GMT. System Locked up - blue screen. Rebooted at 1400 GMT. RTS. VSWR would not reset. (down at 08-04-96 at 0110 GMT).

7/25/96

GRS

1421 GMT. Tried to reset the VSWR light. It would not reset. Left instrument with light flashing.

7/24/96

CAM

1345 GMT. Reset VSWR at 1345 GMT.

7/23/96

CAM

1525 GMT. RF Power controller inop. Breaker tripped during thunderstorms overnight. Reset breaker and RTS O.K. at 1525 GMT.

7/22/96

CAM

1618 GMT. No files being received since system was RTS on the 17th. T/S found fiber optics were never reconnected to computer. Connected fibers, rebooted and RTS at 1618 GMT. All files recovered at next collection.

7/18/96

CAM

1336 GMT. VSWR light flashing. Will not stay reset.

7/17/96

Radian

2000 - 2130 GMT. Radian technician reinstalled phase delay and 3 way splitter. System RTS at 2130 GMT.

6/25/96

CAM

1348 GMT. ALC light ON. Left alone because Radian is on site. Radian powered system down for maintenance. Phase delay and 3 way power splitter sent in for repair.

6/21/96

CAM

1350 GMT. Reset ALC and VSWR. Both continue tp come on.

6/21/96

CAM

1435 GMT. Powered off and removed from network for SDS upgrade. RTS off of network at 1438 GMT.

6/20/96

CAM

1400 GMT. Reset VSWR (several times). RTS at 1404 GMT.

6/19/96

TMG

1418 GMT. Reset VSWR and RTS.

6/18/96

TMG

1351 GMT. Reset ALC and VSWR.

6/17/96

CAM

1430 GMT. Reset ALC and VSWR.

6/14/96

CAM

1432 GM.T Reset ALC and VSWR.

6/13/96

GS/CM

1406 GMT. Daily PM. Reset ALC and VSWR at 1416 GMT. Both lights continue to come on and off. Data looks suspect.

6/12/96

CM/TG

1820 - 1930 GMT. 50 powered down at 1820 GMT. Modified new RASS card as required for our system. Installed card and RTS at 1930 GMT. RASS operational with the 2000 GMT sounding. (SDS indicates data is all 999ís).

6/11/96

CAM

1354 GMT. ALC ON - VSWR flashing. Reset both at 1354 GMT. RTS.

6/11/96

CAM

1418 GMT. Powered down system at 1418 GMT to spray antenna field for weeds.

6/11/96

CAM

2100 - 2135 GMT. Installed new RASS card and loaded RASS only Radar.par but RASS would not sound. Reloaded old Radar.par file and RTS at 2135 GMT. Rich Coulter will research problem and get back in touch with site operations. Meanwhile 50 MHz operating in wind and temperature mode without RASS.

6/10/96

CAM

1412 GMT. VSWR light flashing. Reset at 1412 GMT. RTS.

6/07/96

CAM

1500 GMT. ALC light ON. Output level maxed. Reset ALC at 1500 GMT.

6/06/96

CM/TG

1352 - 1354 GMT. Data looks bad. T/S indicates RF amplifier main circuit breaker tripped. Reset breaker and RTS.

6/05/96

CM/TG

1421 GMT. Reset VSWR and ALC.

6/04/96

CAM

1424 GMT. VSWR light flashing. Reset at 1424 GMT.

6/04/96

CM/TG

1600 - 1737 GMT. Powered down system at 1600 GMT. Re-seated RASS card and inspected for damage. Copied new radar.par file to test in Rass only mode. System still could not find RASS card. Pulled the RASS card out to return to Radian for repair. Copied back the winds only radar.par file. 50 RWP RTS at 1737 GMT in winds only mode.

6/03/96

CAM

1425 - 1506 GMT. No files since 0510 GMT on the first. "RASS DSP ADDR = 0x310 RASS Random Board not at ADDR 0310 - Hit Key to continue". Displayed on monitor. Continued twice but each time system halted again with similar message. Rebooted at 1506 GMT still no files on SDS at 1630 GMT. T/S continues.

6/03/96

CAM

2030 GMT. System stopped again at 1600 GMT. Coulter suspects lightning may have damaged the RASS card. Copied new radar.par file to computer for wind only mode. Rebooted and RTS at 2035 GMT. Waiting to hear from Radian to T/S RASS card.

5/28/96

TMG

1424 GMT. VSWR light flashing. Reset and RTS.

5/27/96

TMG

1535 - 1550 GMT. VSWR light flashing and data looks bad. Reset VSWR light for 10 minutes (VSWR would not reset). RF controller output level needle would peg the gauge and RF power condition would squeal when trying to reset VSWR. Power supply monitoring value increases to approx. 12 Amps when this condition occurs.

5/23/96

GRS

1430 - 1439 GMT. Daily PM. Reset VSWR at 1438 GMT. ALC light lit. T/S continues on VSWR problem.

5/20/96

CAM

1428 - 1450 GMT. Daily PM. VSWR light flashing and will not reset. Powered RF controller off at 1428 GMT. Restored power at 1450 GMT but VSWR light still flashing. RTS as is.

5/16/96

GRS

1450 - 1455 GMT. Daily PM. VSWR flashing. Reset the VSWR at 1453 GMT.

5/14/96

GRS

1514 - 1527 GMT. Daily PM. VSWR flashing. Reset the VSWR at 1514 and again at 1527 GMT. RTS.

5/13/96

MW

1537 - 1552 GMT. The 50 profiler was not sending files to the SDS since Glinda rebooted at 1800 GMT on 05/12/96. Rebooted the 50 computer at 1537 GMT. Collections re-established. Data from 05-12-96 at 1800 GMT to 05-13-96 at 1537 GMT was not collected.

4/18/96

MR

1400 GMT Daily PM. VSWR light flashing. Reset and RTS.

4/17/96

MR

1420 - 1438 GMT. Daily PM. VSWR light flashing. Powered down at 1423 GMT to let controller cool. Powered up at 1430 GMT. RTS at 1438 GMT.

4/16/96

TMG

1351 GMT. Daily PM. VSWR light flashing. Reset at 1351 GMT and light started flashing at 1355 GMT. Powered down RF Power controller at 1400 GMT. Allowed controller to cool. Powered up at 1410 GMT. RTS.

4/15/96

CAM

1400 GMT. Daily PM. VSWR light flashing. Reset at 1400 GMT. Data looks corrupted.

4/12/96

TMG

1925 - 1955 GMT Powered down system to perform PM on computer

4/09/96

CAM

1411 GMT. Wind profiles 999ís since 0611 GMT on the 8th. RF amplifier inop. 60 amp breaker tripped. Reset breaker and RTS at 1411 GMT. (power bump at CF near time of outage).

4/08/96

CAM

1401 GMT. VSWR flashing. Reset and RTS at 1401 GMT.

4/03/96

GRS

1600 GMT. RF amplifier not operating. T/S found 50 Amp breaker tripped on main circuit breaker. Reset breaker and RTS at 1600 GMT.

3/18/96

CAM

Daily checks. VSWR light flashing. Reset and RTS at 1503 GMT. Corrected loose guywire on SE Rass.

3/12/96

DKB

1540 GMT. Powered down system to T/S VSWR. Found nothing. Powered up and tested. The VSWR light reset. Noticed data display looks bad. Made print outs and sent to mentor. RTS at 1600 GMT.

3/11/96

GRS

Daily check. VSWR light flashing and will not reset.

2/13/96

TMG

AC power secured for construction at 1535 GMT. RTS at 2009 GMT.

2/6/96

DKB

Mentor found data collected overnight good. RTS the system at 1500 GMT.

2/5/96

DKB

R/R relay and tested OK. Made print outs and sent to vendor. Ran overnight to test.

1/31/96

DKB

Stopped program to T/S antenna with Radian. Found bad X-Y relay switch but did not have spare. System off line awaiting parts.

1/29/96

TMG

AC power secured for construction at 1453 GMT. RTS at 2110 GMT.

1/12/96

DKB

1430 GMT. Powered down system to T/S with Radian via phone.

1/5/96

DKB

1725 GMT. T/S E-W/ N-S data looks the same. Made print outs of all beam displays and sent them to vendor for interpretation. RTS at 1750 GMT.

12/28/95

CAW/CAM

2015 GMT. T/S North Rass amplifier circuit breaker tripped and will not reset. Found speaker coil bad. R/R and tested OK. RTS at 2210 GMT.

12/27/95

TMG

AC power secured for construction at 1514 GMT.

 12/21/95

TMG

Daily check. VSWR light flashing. Reset and RTS at 1617 GMT.

12/13/95

CAW

Installed the cable and downloaded the .PAR file from the SDS. RTS at 1800 GMT.

12/8/95

MW

* Installed the repaired interface unit and are still awaiting the cable. * New - 6169

10/19/95

TMG

* Installed interface unit and tested bad. Used 915 WP computer to T/S. Found a bad cable in the 50 WP computer and more problems with the interface unit. * Removed the interface unit and sent back to Radian for repair and ordered a new ribbon cable. The system remains Off line. * Installed SN 6169 and removed same

9/26/95

DKB

Mentor has requested T/S on the interface unit. Found the interface unit was not changing the beam direction. * Removed and sent back to Radian for repair. The system is Off line. * Old - 6169

9/18/95

MW

T/S no coms with SDS. Reconfigured computer software and re-booted. RTS at 1522 GMT.

9/15/95

CAM

Powered up and was successful. RTS at 1324 GMT.

9/12/95

GRS

Stopped program to T/S SDS problem. Started back up but the program would not run. Re-booted several times and finally was successful. RTS at 1847 GMT. The program stopped running and the system was left off line until Radian could be contacted.

9/8/95

DKB

Reset VSWR alarm and RTS at 1349 GMT.

9/7/95

DKB

Reset VSWR alarm and RTS at 1325 GMT.

8/31/95

CAM

Powered down for site construction. Powered up and RTS at 2038 GMT.

8/16/95

CAM

Received and * installed the repaired optical disk drive. * New - 2420557

8/15/95

GRS

Reset VSWR alarm and RTS at 1415 GMT.

8/14/95

MW

Reset VSWR alarm and RTS at 1430 GMT.

8/9/95

GRS

Reset VSWR alarm twice with no success. Left VSWR flashing.

8/8/95

DKB

Installed DSP card and tested OK. RTS at 1415 GMT.

7/26/95

DKB

T/S computer error found DSP card bad. Removed card and sent back to Radian for repair. System Off line.

 7/24/95

CAM/MR

Program stopped with " board not at address 310 hit any key to continue" error. 220 breaker also tripped. Reset and restarted with no success. Informed mentor. Left system inop.

7/21/95

DKB

Powered down system to T/S computer problem. Reset circuit board. RTS at 2112 GMT.

7/17/95

CAW

Found program stopped and Rass inop. Powered down system to place the SE Rass speaker upright. Heavy winds overturned it last night.

7/14/95

CAM

Reset VSWR alarm and RTS at 1346 GMT.

7/13/95

CAM

Reset VSWR alarm. RTS at 1335 GMT.

7/12/95

TMG

Reset VSWR alarm and RTS at 1401 GMT. Stopped program and made antenna power measurements using test equipment provided by Radian. Sent the results to Radian for analysis. RTS at 1716 GMT.

7/6/95

GRS

Reset VSWR alarm. RTS at 1434 GMT.

7/4/95

GRS

Reset VSWR alarm but it continued to trip. Left with VSWR flashing.

6/20/95

GRS

Reset VSWR alarm twice.

6/19/95

GRS

Reset VSWR alarm.

6/15/95

CAM

Reset VSWR alarm and made forward/ reflective power measurements. RTS at 1520 GMT.

6/14/95

CAW/CAM

Powered down to check the BSU in the antenna field. Found nothing obvious. RTS at 1525 GMT.

6/13/95

TMG

Reset VSWR alarm 3 times. Informed mentor of the problem. T/S with Radian via phone and made forward and reflective power measurements. RTS at 1830 GMT.

6/5/95

CAM

Reset VSWR alarm.

5/25/95

TMG

T/S system found optical drive bad. * Removed and sent to Radian for repair. Restarted system. Stopped program to check D: size. RTS at 2027 GMT. * Old - 2420557

5/23/95

CAM

T/S system inop with Rass on constantly. Re-boot did not work. Optical drive active constantly. Cleaned and re-initialized with no success. Contacted Radian. System Off line.

5/16/95

GRS

Turned optical disk over to side A.

5/12/95

DKB

Stopped program and tried to copy a data file to the SDS. Was unsuccessful. RTS at 1900 GMT.

 5/11/95

DKB

Powered down system for Main AC power work. RTS at 2109 GMT.

5/1/95

GRS

Program locked up Sat while transferring files to SDS. Re-booted and tested OK. RTS at 1356 GMT.

4/21/95

GRS

Program locked up. Re-initialized optical disk and re-booted. RTS at 1430 GMT.

4/20/95

CAM

Program locked up. Re-booted and RTS at 2121 GMT.

4/16/95

DKB

Turned optical disk over to side B. Rass constantly on. Re-loaded .PAR file and tested bad. T/S found the breaker tripped. Restored and tested OK. RTS at 1230 GMT.

4/7/95

CAM

Turned optical disk over to side A.

4/5/95

DKB

Powered down system to spray antenna field for weeds. RTS at 2145 GMT.

4/4/95

DKB

Made SNR print outs.

3/17/95

CAM

Turned optical disk over to side B.

3/8/95

CAM

Powered off optical disk drive and cleaned. RTS at 1716 GMT.

2/24/95

GRS

Turned optical disk over to side A.

2/10/95

CAM

Powered down to R/R UPS back to the original UPS. RTS at 1515 GMT.

2/3/95

CAM

Turned optical disk over to side B.

1/25/95

DKB

Stopped system to install new .PAR file. System would not run with new file. Loaded old .PAR file and RTS at 2000 GMT.

1/24/95

TMG

Program stopped with "mounting F: and G:" error. Re-booted and RTS at 1440 GMT.

1/23/95

CAM

Program locked up Sun with "F: and G: already in use" error. Powered down system and restarted OK. RTS at 1449 GMT.

1/20/95

CAM

Edited system program to run off Glinda. RTS at 1916 GMT.

1/18/95

CAM

T/S program stopped with "mounting F:, G:" errors. Powered down and restarted OK. RTS at 1445 GMT.

1/11/95

TMG

Stopped program and copied all files from side B of the optical disk to the SDS. Turned the optical disk over and copied all files from side A to the SDS. The SDS did not have sufficient space to copy all files. Informed DMT. DMT made more room and finished copying all files from side A. Del all files from D: and re-initialized optical disk side A. RTS at 1921 GMT.

 1/3/95

GRS

Reset VSWR alarm. The transmitter was in " manual" mode. Placed back into "auto" and RTS at 1545 GMT.

1/2/95

TMG

Display does not look right. Had to reset VSWR alarm on transmitter. Power meter inop. Informed mentor.

12/29/94

CAW

Powered system down to R/R the UPS. RTS at 1430 GMT.

12/28/94

CAM

T/S system inop found UPS bad. Connected system to AC outlets and restarted OK. RTS at 2225 GMT.

12/27/94

TMG

Turned optical disk over to side B.

12/22/94

TMG

T/S Rass inop. Re-boot system and could not find G:. Re-booted again and left with Rass inop. at 1710 GMT.

12/15/94

CAM

Made SNR print outs.

12/7/94

TMG

T/S North Rass speaker inop found amplifier breaker tripped. Reset and tested OK. RTS at 1520 GMT. North Rass speaker went inop later that day. Powered down to T/S. R/R speaker and tested OK. RTS at 2200 GMT.

12/2/94

GRS

T/S Rass on continuously. Powered down and restarted OK. RTS at 1551 GMT.

12/1/94

CAW

Powered down to apply weed control. Powered up and RTS at 2140 GMT.

11/30/94

CAW

Powered down to enter antenna field. Powered up and RTS at 2150 GMT.

11/29/94

DKB

Powered down to defoliate the antenna field. Powered up and RTS at 2150 GMT.

11/16/94

DKB

T/S system inop. Be-boot did not work. Power down and restarted OK. Made SNR print outs. RTS at 1532 GMT.

11/14/94

TMG

T/S system inop since Sun. Re-boot did not work. Powered down everything and restarted OK. RTS at 1544 GMT.

10/31/94

CAM

T/S system inop found program locked up with " device G: already in use". Re-booted and stopped with an " failed to initialize HA" error. Re-booted again OK. RTS at 1440 GMT.

10/27/94

TMG

T/S system inop. Copied new .PAR file over and re-booted. RTS at 1427 GMT.

10/17/94

DKB

T/S system inop found 220 breaker tripped. Reset and tested OK. RTS at 1955 GMT.

10/14/94

GRS

Changed out optical disk.

 10/13/94

GRS

T/S SDS not receiving files. Yates re-booted SDS and tested OK. RTS at 1600 GMT.

10/7/94

DKB

Made SNR print outs.

9/23/94

TMG

Turned optical disk over to side B.

9/21/94

TMG

Reconnected system to SDS and copied files over. everything looks OK. Computer locked up at 1820 GMT. Re-boot and Del .SPD and .SPH files from D:. RTS at 1915 GMT.

9/20/94

CAM

T/S system inop. Re-boot did not work. Copied last CNT, CNW files to SDS and disconnected fiber. Re-booted system and tested OK. Left system disconnected from SDS. RTS at 2120 GMT.

9/19/94

CAW

T/S system inop. Re-booted computer and tested OK. RTS at 1310 GMT. Program locked up again at 1710 GMT. Re-booted computer and Rass started running continuously. Performed hard re-boot and tested bad. Re-loaded .PAR file and tested OK. RTS at 1915 GMT.

9/6/94

DKB

Made SNR print outs.

9/1/94

CAW

T/S system inop found 220 breaker tripped. Reset and tested OK. RTS at 1335 GMT. Powered down final amplifier at 1610 GMT to install new fuse holder.

8/30/94

CAW

System inop with optical disk full. Turned disk over to side A. RTS at 1340 GMT.

8/20/94

GRS

T/S system inop found 220 breaker tripped. Reset and tested OK. RTS at 1338 GMT.

8/12/94

CAM/TMG

T/S system inop found computer locked up since 0010 GMT. Re-booted and tested OK. RTS at 1449 GMT.

8/8/94

TMG

T/S system inop found computer locked up. Re-booted and tested OK. RTS at 1548 GMT.

8/7/94

DKB

Made SNR print outs.

8/6/94

TMG

Turned optical disk over to side B.

7/28/94

DKB

T/S system inop found 220 breaker tripped. Reset and tested OK. RTS at 1715 GMT.

7/25/94

DKB

T/S system inop found computer error had locked the system up. Re-booted and tested OK. RTS at 1344 GMT.

7/21/94

DKB

T/S system inop found 220 breaker tripped. Reset and tested OK. RTS at 1329 GMT.

7/15/94

CAW

R/R optical disk with a re-writeable disk.

 7/13/94

DKB

T/S system inop found 220 breaker tripped. Reset and tested OK. Powered "OFF" system to realign SE Rass speaker. RTS at 1405 GMT.

7/11/94

DKB

T/S no files on SDS. Re-booted and tested OK. RTS at 1927 GMT.

7/6/94

DKB

Made SNR print outs.

6/24/94

CAM

Turned optical disk over to side B.

6/17/94

DKB

Powered "OFF" system to tighten Rass speaker support cables. RTS at 1452 GMT.

6/9/94

CAW

Powered "Off" system to re-install the Rass speakers. RTS at 1605 GMT.

6/6/94

DKB

Made SNR print outs.

6/3/94

CAM/GRS

R/R optical disk.

6/1/94

DN/CAW

Powered system "OFF" to remove Rass speakers. RTS at 1730 GMT. Rass still inop.

5/30/94

DKB

T/S noisy display found 220 breaker to the receivers tripped. Reset and tested OK. RTS at 1330 GMT.

5/25/94

CAW

AC Power fixed and tested. RTS at 1750 GMT.

5/3/94

DKB

T/S Receiver inop found one leg of main AC power bad. Powered down the system and notified electrician.

5/2/94

CAM/GRS

No AC power to trailer outlets. Trailer lights are the only thing working. Turned "OFF" the main breaker and called electrician. Power restored at 1945 GMT. RTS at 1955 GMT. Rass still inop.

4/29/94

CAM/GRS

Found display on "pause". Re-booted and tested OK. RTS at 1440 GMT. Rass still inop.

4/28/94

DN

T/S Rass inop found the speaker transition screws broken off. Powered down Rass amps until repairs on the speakers can be made.

4/18/94

CAM

T/S no Rass. Mentor copied in new .PAR file and tested OK. RTS at 1500 GMT.

4/10/94

DKB

T/S computer inop. Re-booted and tested OK. RTS at 1435 GMT.

 4/4/94

DKB

T/S computer locked up since 4/1 1700. Mentor installed new .PAR file that did not work. Del new and copied old back in. RTS 1430 GMT.